# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ 4 Nonparametric Invariance Pursuit
for neural network architecture hyper-parameter $N,L$. A larger width or depth can be adopted for $\mathcal{F}$ in general. Our specification of width and depth for $\mathcal{F}$ here is for technical purposes, that is, to let $f\in\mathcal{F}$ approximate any $m^{(e,S)}-g$ well for any $g\in\mathcal{G}$.
@@ 4.1 Non-Asymptotic Result for FAIR-NN
Condition 4.5 (Identification for Nonparametric Invariance Pursuit).
@@
$\exists\widetilde{S}\subseteq[d]~{}\text{with}~{}\widetilde{S}\setminus S^{\star}\neq\emptyset\qquad s.t.\qquad\forall e\in\mathcal{E}~{}~{}\mathbb{E}[Y^{(e)}|X_{\widetilde{S}}^{(e)}]\equiv g(X_{\widetilde{S}}^{(e)})~{}~{}\mu^{(e)}\text{-}a.s.~{}~{}\text{with}~{}g:\mathbb{R}^{|S|}\to\mathbb{R},$
in which both set $S^{\star}$ and $\widetilde{S}$ embody the invariant conditional expectation structure, thus more environments are needed in this case to pinpoint $S^{\star}$. Such a minimal identification condition underscores that our proposed FAIR-NN estimator is “sample efficient” regarding the number of environments $|\mathcal{E}|$ required; see the discussions in Section 4.2. Notably, such an identification condition relaxes those employed in approaches using intersections like ICP (Peters et al.,, 2016; Heinze-Deml et al.,, 2018). These approaches require the shifts of conditional distributions for all the $S$ with $\bar{m}^{(S)}\neq m^{\star}$ for identifying $S^{\star}$.
Remark 2.
We claim that 4.5 can be further slightly relaxed given our algorithm searches for the most predictive variable set that preserves the invariance structure. But it is of a technical style and lacks semantic meaning; see discussions in Section A.4.
The following theorem provides an oracle-type inequality for the FAIR-NN estimator in a structure-agnostic manner.
Theorem 4.1 (Oracle-type Inequality for FAIR-NN Least Squares Estimator).
Assume 4.14.5 hold. Then $\gamma^{\star}_{\mathtt{NN}}=\sup_{S\subseteq[d]:\mathsf{b}_{\mathtt{NN}}(S)>0}(\mathsf{b}_{\mathtt{NN}}(S)/\bar{\mathsf{d}}_{\mathtt{NN}}(S))<\infty$, where
@@
Let $C_{1}$ be some large universal constant. Assume further that $|\mathcal{E}|\leq n^{C_{1}}$. Consider the FAIR-NN estimator that solves (3.8) with $\ell(y,u)=\frac{1}{2}(y-v)^{2}$ using $\gamma\geq 8\gamma^{\star}_{\mathtt{NN}}$ and function classes (4.1) with $L,N$ satisfying $NL\leq n^{C_{1}}$ and $N\geq 4$. Then, there exists some constant $\widetilde{C}$ depending on $(d,\sigma_{y},C_{y},b_{m},C_{1})$ such that for any $n\geq 3$,
$\|\widehat{g}-m^{\star}\|_{2}\leq\widetilde{C}(1+\gamma)\left(\delta_{\mathtt{opt}}+\max_{e\in\mathcal{E},S\subseteq[d]}\inf_{h\in[\mathcal{H}_{\mathtt{nn}}(d,L,N,b_{m})]_{S}}\|m^{(e,S)}-h\|_{2,e}+\frac{NL\log^{3/2}n}{\sqrt{n}}\right)=:{\delta}_{\mathtt{NN},1}$
occurs with probability at least $1-\widetilde{C}n^{-100}$. In addition, if $\delta_{\mathtt{NN},1}\leq\widetilde{C}^{-1}\left(1\land s_{\min}\land\{\gamma\inf_{S:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S)\}\right)$, then
$\|\widehat{g}-m^{\star}\|_{2}\leq\widetilde{C}\left(\delta_{\mathtt{opt}}+\max_{e\in\mathcal{E}}\inf_{h\in[\mathcal{H}_{\mathtt{nn}}(d,L,N,b_{m})]_{S^{\star}}}\|m^{\star}-h\|_{2,e}+\frac{NL\log^{3/2}n}{\sqrt{n}}\right)$
occurs with probability at least $1-\widetilde{C}n^{-100}$.
@@
The two quantities $\mathsf{b}_{\mathtt{NN}}(S)$ and $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ can be interpreted as bias mean and bias variance, respectively. It generalizes the quantities defined in Fan et al., (2023) in a nonparametric setting. It matches those up to constant-level factors when $(X^{(e)},Y^{(e)})$ follows a multivariate Gaussian distribution for each $e\in\mathcal{E}$. We refer to $\mathsf{b}_{\mathtt{NN}}(S)$ as bias mean since it exactly characterizes the bias of the least squares estimator in the presence of spurious/endogenous variables like the background color in the thought experiment. In particular, let $\widehat{g}_{\mathtt{LSE}(S)}$ be the least squares estimator that regresses $Y$ on $X_{S}$ using all the data, applying Proposition B.1 gives
@@
We refer to $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ as the bias variance because it measures the variations of bias across environments. Specifically, when $S^{\star}\subseteq S$, the bias in environment $e$ is $(m^{(e,S)}-m^{\star})$, and $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ can be viewed as the variance of the bias concerning the uniform distribution on $\mathcal{E}$ since $\bar{\mathsf{d}}_{\mathtt{NN}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|(m^{(e,S)}-m^{\star})-(\bar{m}^{(S)}-m^{\star})\|_{2,e}^{2}$. We have $\bar{\mathsf{d}}_{\mathtt{NN}}(S^{\star})=0$ by the invariance structure 4.2.
Remark 4 (Main Identification Message).
Theorem 4.1 combines the identification result, which characterizes when it is possible to consistently estimate $m^{\star}$, and the finite-sample estimation error result, which characterizes how accurately we can estimate $m^{\star}$. The main identification message disentangled from the above theorem is that if the minimal heterogeneity condition 4.5 holds, then one can consistently estimate $m^{\star}$ provided $\gamma$ is larger than some threshold $8\gamma^{\star}_{\mathtt{NN}}$ that is independent of $n$.
To present the exact $\ell_{2}$ error rate under a specific nonparametric setup, we first introduce the concept of $(\beta,C)$-smooth function.
@@
One significant advantage of neural networks over traditional nonparametric methods is their intrinsic capability for algorithmic nonparametric regression. This enables them to learn low-dimensional structures with little or no explicit guidance regarding the forms of functions (Bauer & Kohler,, 2019; Schmidt-Hieber,, 2020; Kohler & Langer,, 2021). The following corollary claims that the $\ell_{2}$ estimation error is guarantee for any $n\geq 3$, while for sufficiently large $n$, our proposed estimator can be efficiently adaptive to the hierarchical composition structure of $m^{\star}$ in the same manner. We begin by elucidating the concept of the Hierarchical Composition Model (HCM).
@@
We define function class of hierarchical composition model $\mathcal{H}_{\mathtt{HCM}}(d,l,\mathcal{O},C)$ (Kohler & Langer,, 2021) with $l,d\in\mathbb{N}^{+}$, $C\in\mathbb{R}^{+}$, and $\mathcal{O}$, a subset of $[1,\infty)\times\mathbb{N}^{+}$ satisfying $\sup_{(\beta,t)\in\mathcal{O}}(\beta\lor t)<\infty$, in a recursive way as follows. Let $\mathcal{H}_{\mathtt{HCM}}(d,0,\mathcal{O},C)=\{h(x)=x_{j},j\in[d]\}$, and
@@
Following Kohler & Langer, (2021), we assume all the compositions are at least Lipschitz functions to simplify the presentation. The minimax optimal $L_{2}$ estimation risk over $\mathcal{H}(d,l,\mathcal{O},C_{h})$ is $n^{-\alpha^{\star}/(2\alpha^{\star}+1)}$, where $\alpha^{\star}=\min_{(\beta,t)\in\mathcal{O}}(\beta/t)$ is the smallest dimensionality-adjusted degree of smoothness that represents the hardest component in the hierarchical composition (Fan et al.,, 2022; Schmidt-Hieber,, 2020). Now we are ready to present the optimal rate for FAIR-NN when all the associations $m^{(e,S)}$ lies in potentially larger HCM, and $m^{\star}$ resides within smaller HCM.
Condition 4.6 (Function Complexity and Neural Network Architecture).
@@
(1) $m^{(e,S)}\in\mathcal{H}_{\mathtt{HCM}}(|S|,l,\mathcal{O},C_{h})$ for any $e\in\mathcal{E}$ and $S\subseteq[d]$ with $\alpha_{0}=\inf_{(\beta,t)\in\mathcal{O}}(\beta/t)$.
(2) $m^{\star}\in\mathcal{H}_{\mathtt{HCM}}(|S^{\star}|,l,\mathcal{O}^{\star},C_{h})$ with $\alpha^{\star}=\inf_{(\beta,t)\in\mathcal{O}^{\star}}(\beta/t)$.
(3) $\delta_{\mathtt{opt}}=o(1)$, we choose $N,L$ satisfying $LN\asymp\{n(\log n)^{8\alpha^{\star}-3}\}^{\frac{1}{2(2\alpha^{\star}+1)}}$ and $(\log n)/(N\land L)=o(1)$
(4) $\max\{C_{1},d,l,C_{h},\sigma_{y},C_{y},b_{x},b_{m},\sup_{(\beta,t)\in\mathcal{O}}(\beta\lor t),\sup_{(\beta,t)\in\mathcal{O}^{\star}}(\beta\lor t)\}\leq C_{2}$ for some constant $C_{2}>0$.
Corollary 4.2 (Optimal Rate for FAIR-NN).
@@
Under the setting of Theorem 4.1, assume further that 4.6 holds. Then, for any $n\geq 3$, with probability at least $1-\widetilde{C}n^{-100}$, the following holds
$\|\widehat{g}-m^{\star}\|_{2}\leq\widetilde{C}\left[\delta_{\mathtt{opt}}+\left(\frac{\log^{7}n}{n}\right)^{-\frac{\alpha^{\star}}{2\alpha^{\star}+1}}\right]+1_{\{n<n_{0}\}}\widetilde{C}\gamma\left[\delta_{\mathtt{opt}}+\left(\frac{\log^{7}n}{n}\right)^{-\frac{\alpha_{0}}{2\alpha^{\star}+1}}\right],$ (4.3)
@@
where $n_{0}$ is a constant dependent on $(C_{2},\delta_{\mathtt{opt}},\gamma,s_{\min},\inf_{S:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S))$, and $\widetilde{C}$ is a constant dependent on $C_{2}$ but independent of $\gamma,\delta_{\mathtt{opt}}$ and $n$.
The error bound (4.3) is applicable for any $n\geq 3$. Notably, the error bound will not inflate if the signal of important variables $s_{\min}$ and the heterogeneity signal $\inf_{S\subseteq[d]:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ is small. Though the general error bound scales linearly with $\gamma$, the estimator we propose is not vulnerable to “weak spurious” variables, for example, $x_{j}$ where $\|m^{(e,S^{\star}\cup\{j\})}-m^{\star}\|_{2,e}\leq\epsilon$ for any $e\in\mathcal{E}$, provided all the ratio of the bias $\mathsf{b}_{\mathtt{NN}}(S)$ to heterogeneity $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ gets controlled. When $n$ larger than some constant $n_{0}$ that may inflate as the signals $(s_{\min},\inf_{S:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S))$ goes to zero, the second term in (4.3) disappears and the rate is (1) independent of $\gamma$ and $\alpha_{0}$; and (2) only dependent on the $\alpha^{\star}$.
@@
Utilizing neural network class for both predictor and discriminator function classes allows the estimator to adapt to the invariant regression function $m^{\star}$ efficiently from two crucial perspectives. Firstly, similar to using neural networks in nonparametric regression (Bauer & Kohler,, 2019; Kohler & Langer,, 2021; Schmidt-Hieber,, 2020), implementing the predictor function using neural network endows the estimator with the capability of being adaptive to the low-dimensional hierarchical structure algorithmically. Secondly, the choice of model parameter $(N,L)$, and the convergence rate only depends on $m^{\star}$. The (spurious) conditional expectations $m^{(e,S)}$ can be much more complex than $m^{\star}$. Notably, this complexity will not affect the convergence rate. This can be credited to the scalability of neural networks used as discriminators, i.e., their adaptivity capability in the role of regularization in FAIR.
@@
At the same time, though we have to choose a hyper-parameter $\gamma$ larger than a certain threshold to attain such a rate, the convergence rate is independent of $\gamma$. This implies that when there is a lot of data available, we do not need to tune the hyper-parameter $\gamma$ for optimal performance carefully. Instead, we can choose some conservative (large) $\gamma$ such that the lower bound requirement $\gamma\geq 8\gamma^{\star}_{\mathtt{NN}}$ is guaranteed.
The statistical rate of convergence (4.3) is minimax optimal (up to logarithmic factors). It is noteworthy that this rate matches the case where we have specific prior cause-effect knowledge such that all the variables leading to a biased estimation ($\bar{m}^{(S\cup S^{\star})}\neq m^{\star}$) are all screened out before fitting the model. We impose only one minimal 4.5 besides invariance assumption and standard assumptions used in nonparametric regression to achieve this. The discussion in the following section asserts that this assumption can be easily satisfied with $S^{\star}$ coinciding with the direct causes of $Y$ in the setting of SCM with interventions on $X$ even when $|\mathcal{E}|=2$, which highlights the capability of our proposed estimator to unveil the causality from multi-source observational data in a complete algorithmic manner, which can learn the causal mechanism between the target variable and its direct causes with no explicit supervision of cause-effect relationship and knowledge of intervention.
4.2 The Semantics of $S^{\star}$ in the SCM with Interventions Setting

### v2

@@ 2.2 Proposed FAIR-NN Least Squares Estimator
for neural network architecture hyper-parameters $N,L$ and truncation parameter $B=C_{0}$. Here $B$ can be larger than $C_{0}$ but should satisfies $B=O(1)$. A larger width, depth, and truncation parameter can also be adopted for $\mathcal{F}$. Our specification of $(N,L,B)$ for $\mathcal{F}$ here is for technical purposes, that is, any $m^{(e,S)}-g$ for $g\in\mathcal{G}$ can be well approximated by some $f\in\mathcal{F}$.
@@ 2.3 Non-Asymptotic Result for FAIR-NN
Condition 2 (Identification for Nonparametric Causality Pursuit).
@@
$\exists\widetilde{S}\subseteq[d]~{}\text{with}~{}\widetilde{S}\setminus S^{\star}\neq\emptyset\qquad s.t.\qquad\forall e\in\mathcal{E}~{}~{}\mathbb{E}[Y^{(e)}|X_{\widetilde{S}}^{(e)}]\equiv g(X_{\widetilde{S}}^{(e)})~{}~{}\mu^{(e)}\text{-}a.s.~{}~{}\text{for some}~{}g:\mathbb{R}^{|S|}\to\mathbb{R},$
in which both set $S^{\star}$ and $\widetilde{S}$ embody the invariant conditional expectation structure, thus more environments are needed in this case to pinpoint $S^{\star}$. Such a minimal identification condition underscores that our proposed FAIR-NN estimator is “sample efficient” regarding the number of environments $|\mathcal{E}|$ required; see the discussions in Section 3. Notably, such an identification condition relaxes those employed in approaches using intersections like ICP (Peters et al.,, 2016; Heinze-Deml et al.,, 2018). These approaches require the shifts of conditional distributions for all the $S$ with $\bar{m}^{(S)}\neq m^{\star}$ for identifying $S^{\star}$.
Remark 2 (Relaxing 2).
We claim that 2 can be slightly relaxed given our algorithm searches for the most predictive variable set that preserves the invariance structure. But it is of a technical style and lacks semantic meaning; see discussions in Section A.4.
The following theorem provides an oracle-type inequality for the FAIR-NN least squares estimator in a structure-agnostic manner. The first term is the maximum approximation bias of neural networks across environments and the second term is related to the complexity of the neural networks used in the fitting. It implies that when the FAIR-NN penalty parameter $\gamma$ is large enough, all endogenous spurious variables can be surely screened Fan & Lv, (2008) when $n$ is large enough, thus $m^{\star}$ can be estimated as well as if the invariant quasi-causal set of variables $S^{\star}$ is known. In addition, the theorem quantifies the amount of penalty needed, which is related to the signal-to-noise ratio of the problem.
Theorem 1 (Oracle-type Inequality for FAIR-NN Least Squares Estimator).
Assume 1 and 2 hold. Then $\gamma^{\star}_{\mathtt{NN}}=\sup_{S\subseteq[d]:\mathsf{b}_{\mathtt{NN}}(S)>0}(\mathsf{b}_{\mathtt{NN}}(S)/\bar{\mathsf{d}}_{\mathtt{NN}}(S))<\infty$, where
@@
Consider the estimator that solves (2.1) using $\gamma\geq 8\gamma^{\star}_{\mathtt{NN}}$ and function classes (2.3) with $L,N$ satisfying $NL\leq n$ and $N\geq 4$. Then, there exists some constant $\widetilde{C}$ depending on $(d,C_{0})$ such that for any $n\geq 3$,
$\frac{\|\widehat{g}-m^{\star}\|_{2}}{\widetilde{C}}\leq\max_{e\in\mathcal{E}}\inf_{h\in\mathcal{G}_{S^{\star}}}\|m^{\star}-h\|_{2,e}+\frac{NL\log^{3/2}n}{\sqrt{n}}+1_{\{{\delta}_{\mathtt{NN},1}>\mathsf{s}\}}\cdot\left(\gamma{\delta}_{\mathtt{NN},1}\right)$
occurs with probability at least $1-\widetilde{C}n^{-100}$. Here $\delta_{\mathtt{NN},1}=\max_{e\in\mathcal{E},S\subseteq[d]}\inf_{h\in\mathcal{G}_{S}}\|m^{(e,S)}-h\|_{2,e}+\frac{NL\log^{3/2}n}{\sqrt{n}}$ and $\mathsf{s}=\widetilde{C}^{-1}[1\land s_{\min}\land\{\gamma\inf_{S:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S)\}]/(1+\gamma)$, where $s_{\min}$ is defined in 1(5).
As our result is non-asymptotic, for a given $n$, we may not be able to eliminate all endogenous spurious variables. The third term in Theorem 1 reflects this when the signal is not sufficiently large. It is more explicitly given in Corollary 1.
@@
We refer to $\mathsf{b}_{\mathtt{NN}}(S)$ as bias mean since it exactly characterizes the bias of the least squares estimator in the presence of endogenous spurious variables like the background color in the thought experiment. In particular, letting $\widehat{g}_{\mathtt{LSE}(S)}$ be the least squares estimator that regresses $Y$ on $X_{S}$ using all the data, namely, the FAIR-NN estimator with $\gamma=0$, Proposition 6 implies
@@
We refer to $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ as the bias variance because it measures the variations of bias across environments. Specifically, when $S^{\star}\subseteq S$, the bias in environment $e$ is $(m^{(e,S)}-m^{\star})$, and $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ can be viewed as the variance of the bias concerning the uniform distribution on $\mathcal{E}$ since $\bar{\mathsf{d}}_{\mathtt{NN}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|(m^{(e,S)}-m^{\star})-(\bar{m}^{(S)}-m^{\star})\|_{2,e}^{2}$. We have $\bar{\mathsf{d}}_{\mathtt{NN}}(S^{\star})=0$ by the invariance structure in 1(b).
Remark 4 (Identification).
Theorem 1 combines the identification result, which characterizes when it is possible to consistently estimate $m^{\star}$, and the finite-sample estimation error result, which characterizes how accurately we can estimate $m^{\star}$. The main identification message disentangled from the above theorem is that if the minimal heterogeneity condition 2 holds, then one can consistently estimate $m^{\star}$ provided $\gamma$ is larger than some threshold $8\gamma^{\star}_{\mathtt{NN}}$ that is independent of $n$.
2.4 Adapting to the Low-dimensional Structures Algorithmically
To present the explicit $L_{2}$ error rate under a specific nonparametric setup, we first introduce the concept of $(\beta,C)$-smooth function.
@@ 2.4 Adapting to the Low-dimensional Structures Algorithmically
One significant advantage of neural networks over traditional nonparametric methods is their intrinsic capability for algorithmic nonparametric regression. This enables them to learn low-dimensional structures with little or no explicit guidance regarding the forms of functions (Bauer & Kohler,, 2019; Schmidt-Hieber,, 2020; Kohler & Langer,, 2021; Fan & Gu,, 2024). We begin by elucidating the concept of the Hierarchical Composition Model (HCM), which is basically the compositions of $t$-variate functions with $(\beta,C)$-smooth $l$ times for $(t,\beta)$ in a certain set $\mathcal{O}$.
@@
We define function class of hierarchical composition model $\mathcal{H}_{\mathtt{HCM}}(d,l,\mathcal{O},C)$ (Kohler & Langer,, 2021) with $l,d\in\mathbb{N}^{+}$, $C\in\mathbb{R}^{+}$, and $\mathcal{O}$, a subset of $[1,\infty)\times\mathbb{N}^{+}$, in a recursive way as follows. Let $\mathcal{H}_{\mathtt{HCM}}(d,0,\mathcal{O},C)=\{h(x)=x_{j},j\in[d]\}$, and for each $l\geq 1$,
@@
Following Kohler & Langer, (2021), we assume all the compositions are at least Lipschitz functions to simplify the presentation. The minimax optimal $L_{2}$ estimation risk over $\mathcal{H}(d,l,\mathcal{O},C_{h})$ is $n^{-\alpha^{\star}/(2\alpha^{\star}+1)}$, where $\alpha^{\star}=\min_{(\beta,t)\in\mathcal{O}}(\beta/t)$ is the smallest dimensionality-adjusted degree of smoothness that represents the hardest component in the composition. For example, if $m^{\star}(x)=f_{1}(x_{1})+f_{2}(f_{3}(x_{2},x_{3}),f_{4}(x_{4},x_{5}))+f_{5}(x_{1},x_{3},x_{5})$ and all functions have a bounded second derivative, then the hardest component is the last one, and the dimensionality-adjusted degree of smoothness is $\alpha^{*}=2/3$.
Condition 3 (Function Complexity and Neural Network Architecture).
@@
(a) $m^{(e,S)}\in\mathcal{H}_{\mathtt{HCM}}(|S|,l,\mathcal{O},C_{h})$ for any $e\in\mathcal{E}$ and $S\subseteq[d]$ with $\alpha_{0}=\inf_{(\beta,t)\in\mathcal{O}}(\beta/t)$.
(b) $m^{\star}\in\mathcal{H}_{\mathtt{HCM}}(|S^{\star}|,l,\mathcal{O}^{\star},C_{h})$ with $\alpha^{\star}=\inf_{(\beta,t)\in\mathcal{O}^{\star}}(\beta/t)$.
(c) We choose $N,L$ satisfying $LN\asymp\{n(\log n)^{8\alpha^{\star}-3}\}^{\frac{1}{2(2\alpha^{\star}+1)}}$ and $(\log n)/(N\land L)=o(1)$
(d) $\max\{C_{0},d,l,C_{h},\sup_{(\beta,t)\in\mathcal{O}}(\beta\lor t),\sup_{(\beta,t)\in\mathcal{O}^{\star}}(\beta\lor t)\}\leq C_{1}$ for some constant $C_{1}>1$.
Corollary 1 (Optimal Rate for FAIR-NN).
@@
Under the setting of Theorem 1, assume further that 3 holds. Then, for any $n\geq 3$, with probability at least $1-\widetilde{C}n^{-100}$, the following holds
$\frac{\|\widehat{g}-m^{\star}\|_{2}}{\widetilde{C}\log^{7}(n)}\leq n^{-\alpha^{\star}/(2\alpha^{\star}+1)}+1_{\{n<n_{0}\}}\gamma\cdot n^{-\alpha_{0}/(2\alpha^{\star}+1)},$ (2.5)
@@
where $n_{0}$ depends on $(C_{1},\gamma,s_{\min},\inf_{S:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S))$, and $\widetilde{C}$ is a constant dependent only on $C_{1}$.
@@
From Corollary 1, we can get (up to logarithmic factors) minimax convergence rate $n^{-\alpha^{\star}/(2\alpha^{\star}+1)}$, which is independent of both $\alpha_{0}$ and $\gamma$, when $n$ is larger than some constant $n_{0}$. Utilizing neural networks in predictor and discriminator function classes allows the estimator to adapt to the invariant regression function $m^{\star}$ efficiently from two crucial perspectives. Firstly, similar to using neural networks in nonparametric regression (Schmidt-Hieber,, 2020; Kohler & Langer,, 2021; Fan & Gu,, 2024), adopting neural networks in $\mathcal{G}$ endows the estimator with the capability of being adaptive to the low-dimensional hierarchical structure algorithmically. Secondly, the choice of model parameter $(N,L)$, and the convergence rate depends only on $m^{\star}$. The (spurious) conditional expectations $m^{(e,S)}$ can be much more complex than $m^{\star}$. Notably, this complexity will not affect the convergence rate. This can be credited to the scalability of neural networks used as discriminators, i.e., their adaptivity capability in the regularization part of FAIR.
Remark 5 (Guaranteed for All $n$).
The error bound (2.5) is applicable for any $n\geq 3$, even when it selects the wrong variables. Notably, the error bound will not inflate if the invariant signal $s_{\min}$ and the heterogeneity signal $\inf_{S\subseteq[d]:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ is small. Though the error bound scales linearly with $\gamma$, the estimator we propose is not vulnerable to “weak spurious” variables, e.g., $x_{j}$ with $\sup_{e\in\mathcal{E}}\|m^{(e,S^{\star}\cup\{j\})}-m^{\star}\|_{2,e}\leq\epsilon$, provided all the ratio of the bias $\mathsf{b}_{\mathtt{NN}}(S)$ to heterogeneity $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ gets controlled.
Remark 6 (Choice of the Hyper-parameter $\gamma$).
@@
Though we have to choose a hyper-parameter $\gamma$ larger than a certain threshold to attain such a rate, the convergence rate is independent of $\gamma$. This implies that when the sample size $n$ is large, we do not need to tune the hyper-parameter $\gamma$ for optimal performance. Instead, we can choose some conservative (large) $\gamma$ such that the lower bound $\gamma\geq 8\gamma^{\star}_{\mathtt{NN}}$ is guaranteed.
3 Nonparametric Invariance Pursuit under SCMs